# Improving the Distributional Alignment of LLMs using Supervision

### v2

@@ Abstract
The ability to accurately align LLMs with human population groups on subjective questions would have great value. In this work, we show that use of simple supervision can greatly improve language model alignment with diverse population groups more consistently, as measured over three datasets spanning various topics. Beyond evaluating average alignment, we also report how alignment varies across specific groups. Our broad findings provide insights into the distributional alignment of LLMs with diverse population groups. By conducting evaluation over many LLMs and prompting strategies, along with open-sourcing our work, we provide a benchmark to stimulate future research. [^1] Our data and code will be available at https://github.com/GauriKambhatla/supervised-llm-alignment
@@ 1 Introduction
As LLMs are increasingly incorporated into real-world systems, it is more and more important to evaluate their alignment with human values and opinions. A growing body of work seeks to align LLMs with particular human population groups, with the goal of simulating their opinions or behaviors, by steering LLM generations towards those groups Aher et al. (2023); Sun et al. (2024); Dillion et al. (2023). One common method to do so is persona or sociodemographic (SD) prompting, which refers to defining a persona with demographic or description variables within a prompt, with the goal of modeling the behavior of that group using an LLM Hu and Collier (2024); Beck et al. (2024).
@@
Recent work in NLP has sought to evaluate SD prompting by studying its performance on large surveys and polls, which contain many subjective questions, along with hundreds of thousands of human responses Santurkar et al. (2023); Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). However, evaluation of SD prompting has typically suffered from one or more key shortcomings. First, many studies have assumed a single majority answer to each question for each sociodemographic group in modeling and/or evaluation Hwang et al. (2023); Hu and Collier (2024); Sun et al. (2025a); Mukherjee et al. (2024); Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). The fallacy in such a framing is clear: because members of a given group do not all share the same beliefs, accurate modeling and evaluation must incorporate intra-group disagreement. Prompting for a distribution, rather than a representative response, could potentially also make a model less susceptible to caricature and stereotyping Cheng et al. (2023a); Wang et al. (2025a); Cheng et al. (2023b), better aligning with the goals of pluralistic alignment Sorensen et al. (2024). However, Prior studies modeling distributional beliefs have also tended to evaluate only a single method of extracting an LLM distribution Santurkar et al. (2023); Sun et al. (2024), or evaluate LLM distributions prompted for specific demographics in a demographic-agnostic manner Beck et al. (2024).
@@
In this work, we evaluate SD prompting in a distributional manner across three survey datasets using a variety of methods to extract distributions from LLMs. We focus on methods that work most broadly, supporting use with open-source, open-model, and API-access LLMs. We apply supervised calibration to better align LLM-generated distributions with human response distributions, with the intuition that LLM distributions might be directionally correct, but simply uncalibrated; e.g., LLM distributions might exaggerate differences between different groups. We use supervised calibration to mean learning a transformation to a distribution to be more similar to human response distributions, using human responses as supervised labels (described in detail in Section 3.6). We use alignment to mean the degree to which opinions align between humans and LLMs (see Section 3.4).
@@
Concretely, we investigate distributional opinion alignment via the following research questions: (RQ 1) Does SD prompting generate distributions that are more aligned with human opinion? (RQ 2) Can we calibrate LLM-generated distributions to be more aligned with humans through supervised data? Is this consistent across models, datasets, and distribution elicitation methods? (RQ 3) Can generated distributions be more easily aligned with some SD groups over others? What is the effect of calibration for individual groups? (RQ 4) Does calibration work with fewer supervised training examples? And (RQ 5) What are the effects of post-training (doing fine-tuning and reinforcement learning) on LLM-generated distributions?
@@
We evaluate the alignment of LLM-generated distributions to human opinion across 3 distribution elicitation methods and 15 models of varying degrees of openness (open-source, open-model, API-access only), size, pre-training data, and post-training methods. We first show that baseline SD prompting techniques do not lead to consistent improvements in alignment across settings (RQ1). In contrast, our approach to supervised calibration of LLM-generated distributions improves alignment by 16.3% on average across prompt methods, elicitation methods, datasets, and LLMs (RQ2), though it improves for some SD groups and degrades for others (RQ3). Moreover, as few as 1-10 gold examples can be used to calibrate distributions (RQ4). In regard to LLM post-training (RQ5), calibration results in more consistent behavior across various post-training methods, both with and without SD prompting; without calibration, alignment is much more dependent on both the dataset and the post-training method used.
We conclude with a discussion of the methods and models that generate the distributions most aligned with human groups on subjective questions. Caution is warranted given varying success in aligning to different groups as well as recognition that such efforts risk essentializing sociodemographic categories if not applied with care.
@@ 2 Related Work
LLMs are increasingly used to simulate social human groups in a variety of settings, such as psychology, sociology, and economic studies Aher et al. (2023); Dillion et al. (2023), survey development Rothschild et al. (2024); Jansen et al. (2023), NLP annotation Nasution and Onan (2024); Hashemi et al. (2024); Movva et al. (2024), and multi-agent interactive simulation Hou et al. (2025); Zhou et al. (2024).
Many methods have been proposed to better align LLMs with human values and opinions Ma et al. (2024). For example, persona or sociodemographic (SD) prompting defines a persona with demographic or description variables within a prompt to simulate or model groups Hu and Collier (2024); Beck et al. (2024). Other methods include: using knowledge of a user’s previous opinions to model their future opinions Hwang et al. (2023); Gordon et al. (2022), using psychological scaffolds to improve model rationalization Joshi et al. (2025), self-alignment via the simulation of social scenes Pang et al. (2024), extracting persona vectors to mitigate certain character traits Chen et al. (2025), creating more descriptive value profiles Sorensen et al. (2025), using in-context learning Choenni and Shutova (2025), and fine-tuning LLMs followed by collection of a small amount of human data Krsteski et al. (2025). Evaluation frameworks have also been proposed for assessing alignment and social simulation Kang et al. (2025); Shi et al. (2025); Lutz et al. (2025).
@@
Various limitations have been reported, such as LLMs being unreliable as personalized annotators Dong et al. (2024), unable to use different languages to steer generation towards groups Kwok et al. (2024), biased towards some groups Sun et al. (2025b), lacking stability and steerability under prompt variations Khan et al. (2025), and lacking the reasoning depth to simulate humans in social experiments Gao et al. (2025). Prior work evaluating SD prompting has also reported mixed results in accurately simulating social groups Beck et al. (2024); Santurkar et al. (2023); Hu and Collier (2024); Kaiser et al. (2025). While some works show that SD prompting can improve performance, results vary greatly by prompt, model, and task Beck et al. (2024); Mukherjee et al. (2024); Bisbee et al. (2024).
Other work warns against such methods due to misportrayal, othering, and exoticization of identities Cheng et al. (2023a); Wang et al. (2025a); Cheng et al. (2023b); Kitadai et al. (2024). Simply incorporating demographic features into prompts does not always enable LLMs to adopt the desired perspectives of specified demographic groups Sun et al. (2025a). Furthermore, a vast body of work has established biases in language models that may be perpetuated with SD prompting Nadeem et al. (2021); Gallegos et al. (2024); Gupta et al. (2024).
Probabilistic reasoning with LLMs
Prior work in NLP has studied the use of LLMs for eliciting and working with uncertainty and probability distributions. Paruchuri et al. (2024) study whether LLMs can reason about probability distributions, finding that LLMs are able to answer questions about distributions, though performance varies based on the type of distribution. Wang et al. (2025b) develop a model and framework for modeling and evaluating fine-grained probability estimation of propositions conditioned on context, and show that their model captures some amount of human uncertainty. Prior work has also found that several LLMs are capable few-shot regressors, often outperforming traditional supervised methods Vacareanu et al. (2024), and capable fine-tuned regressors Song et al. (2024).
Distribution prediction in the real world is a complex task, and can rely on both uncertainty quantification and nuanced language understanding Wang et al. (2025b); Vedula et al. (2025). This motivates using LLMs for this task, and prior work has developed methods in various domains for doing LLM-based distribution prediction: combining LLM-generated distributions to better predict human annotator distributions Hashemi et al. (2024), developing a quantile regression approach for price prediction Vedula et al. (2025), and mapping LLM token predictions to continuous distributions for probabilistic forecasting Gruver et al. (2023).
@@
Recent works have also advocated for a shift to pluralistic alignment due to the ecological fallacy of assuming all members of a group share the same values, beliefs, or opinions Gordon et al. (2022); Sorensen et al. (2024); Feng et al. (2024); Kirk et al. (2024); Xie et al. (2025). Closer to our work, prior works have studied aligning LLM distributions with human response distributions on subjective questions. We draw inspiration from Santurkar et al. (2023), though we evaluate multiple methods of extracting distributions from LLMs. Dominguez-Olmedo et al. (2024) study systematic biases with LLM-generated distributions and how this differs from human response distributions for surveys. Meister et al. (2025) study different distribution elicitation methods for subjective questions. Cao et al. (2025) fine-tune with a divergence-based loss function. Our work introduces the notion of supervised calibration in this space, and studies the effect of alignment methods on individual groups.
@@ 3.1 Datasets
We consider three survey datasets: Welcome Global Monitor 2018 (WGM), OpinionQA (OQA) Santurkar et al. (2023), and the World Values Survey (WVS). OQA and WVS have been used to study alignment in prior work Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). These three datasets cover diverse subjective topics, including global perceptions of science and public health (WGM), US public opinion (OQA), and various moral opinions and values (WVS). We include all WGM ordinal questions and two questions per category from OQA and WVS, yielding 92 questions across WGM (14), OQA (38), and WVS (40) and a total of 4,500 human response distributions over all datasets, questions, and demographics.
@@ 3.2 Distribution elicitation
From ground truth human response data, we infer reference probability distributions over answer choices for each question by relative frequency, specific to each SD group. To generate distributions via LLM, we apply (with some modifications) methods inspired from model confidence literature. We investigate techniques that work most broadly, supporting use with open-source, open-model, and API-access LLMs. That said, Appendix A.2 reports a smaller scale study with log probabilities. Appendix B shows prompts used for all methods.
Verbalized. This is directly prompting the model to output a probability distribution Geng et al. (2024); Tian et al. (2023). For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a distribution over the answer choices, such as [0.7, 0.2, 0.05, 0.05]. We sample $n=3$ times and average for variance reduction. If a model ever does not generate a distribution as requested, we simply discard that output. We also renormalize generated distributions that do not sum to 1.
Self-random. We prompt the model to output a single answer choice, sampling $n$ times Xiong et al. (2024), with a temperature of 0.7. We then create a distribution over the $n$ responses. For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a single number that correlates with an answer choice, such as 4 for a lot or 1 for not at all. Note that this method estimates the log probabilities for large $n$ Tian et al. (2023). In our experiments, we sample $n=5$ times, following prior work Xiong et al. (2024).
Paraphrase. We also prompt the model to output a single answer choice but use different paraphrases of the prompt instead of sampling responses for the same prompt Xiong et al. (2024). We use $n=5$ paraphrases for each prompt.
@@ 3.3 Prompts
We use two categories of prompts, with three variations. For each, we prompt with the 3 different elicitation methods. See Appendix B for the prompts.
@@
Prompt variations. We study three prompt variations, standard prompts, few shot prompts (Section 6.2), and chain of thought (CoT) prompts Kojima et al. (2022); Wei et al. (2022) (Section 4.3). See Appendix Table 16 for the variations.
@@ 3.4 Metric: Opinion Alignment
where $WD$ refers to Wasserstein distance, or earth-mover’s distance, and N1N-1 is a normalization factor. Santurkar et al. (2023) note this metric accounts for the ordinal nature of the survey questions, as opposed to other distribution divergence metrics like Kullback-Liebler or Jensen-Shannon. The opinion alignment metric ranges in $[0,1]$, but in our results we show the value as a percentage.
@@ 3.5 Models
We evaluate 15 models ranging in openness (open-source, open-weight, black box), size, modality, and post-training. Model families include Claude, Llama, Mistral, OLMo-2, and Qwen. Our main results report on only the most powerful model in each family. Others are reported in Appendix A.1.
@@
Table 1: Opinion alignment before and after calibration for each dataset, LLM, and elicitation method. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for each elicitation method: paraphrase ($P$), self-random ($S$), and verbalized ($V$). Bolded values are significant between each pair (calculated via paired t-test and Bonferroni correction). The best performance for each model is highlighted in yellow. The best averages across models and datasets are highlighted in blue. Results for all LLMs are shown in Appendix Table 7. Adding SD information does not consistently improve alignment, but calibration more consistently improves alignment on average and in most settings.
@@ 3.6 Calibration
We apply supervised regression to transform LLM-generated distributions to more closely align them with the ground truth human distributions. For example, for the question, How much do you trust doctors and nurses? A lot, some, not much, or not at all?, the LLM might generate a distribution such as $[0.6,0.2,0.2,0]$, while the ground truth human response distribution based on the survey question might be $[0.25,0.5,0.1,0.15]$. We seek to learn a transformation from the LLM-generated distributions to the human distributions that might be applied to all LLM-generated distributions. More formally, for a regression model $R$ and LLM distribution $\mathbf{D}$ parameterized by the values for each answer choice $D_{a},D_{b},D_{c},...,D_{k}]$ (where $k$ is the number of answer choices), we learn a regression such that each value is transformed using supervision from ground truth values for each answer choice $G_{a},G_{b},G_{c},...,G_{k}]$. We then re-normalize the transformed distribution.
@@
Regression model. We learn a regression model for each dataset-LLM-elicitation-prompt setting (270 regression models for 3 datasets, 15 models, 3 elicitation methods, and 2 prompt categories). Hyperparameter tuning and model selection are used to choose between linear regression (including lasso and ridge) and random forest regression for each regressor. We use the scikit-learn Pedregosa et al. (2011) implementation for each regressor.
Training. The total number of examples for each regression model is $C$ x $Q$ x $A$, where $C$ is the class within each sociodemographic group (e.g., female and male for sex), $Q$ is the question (e.g., how much do you trust doctors?), and $A$ is the number of answer choices per question. These values are different for each setting (see Table 13 for the sizes of $Q$ and $C$ per dataset). We were unable to extract distributions for all questions from some of the smaller LLMs (see Table 17), so the number of training examples differs for them. We split the LLM distributions into train, development, and test sets (80%-10%-10%).
For minimal supervision (Section 6), we train on a random sample of $n$ examples ($n$ = [1, 5, 10, 50, 100, 200, 500, 1000, 1200], where each example is the full distribution not split by answer choices) from the training set. We repeat this procedure 10 times for different random splits and average.
Evaluation. We optimize the regression models with Mean Squared Error (MSE). Since we split distributions during training, we normalize output values so that they sum to 1 before calculating opinion alignment with the ground truth distribution. For example, if the output of $R$ for the distribution $[0.6,0.2,0.2,0]$ is $[0.3,0.7,0.1,0.1]$, we would normalize these values to $[0.25,0.59,0.08,0.08]$.

### v3

@@ Abstract
The ability to accurately align LLMs with population groups on subjective questions would have great value. In this work, we show that simple supervision can more consistently improve language model alignment with diverse population groups, as measured across three datasets spanning various topics. Beyond evaluating average alignment, we also report how alignment varies across specific groups. Our broad findings provide insights into the distributional alignment of LLMs with diverse populations. By conducting evaluation over many LLMs and prompting strategies, we provide a benchmark to stimulate future research. [^1] Our data and code will be available upon publication.
Improving the Distributional Alignment of LLMs using Supervision
Gauri Kambhatla, Sanjana Gautam, Angela Zhang, Alex Liu, Ravi Srinivasan, Junyi Jessy Li, Matthew Lease The University of Texas at Austin gkambhat@utexas.edu
@@ 1 Introduction
As LLMs are increasingly incorporated into real-world systems, it is more and more important to evaluate their alignment with human values and opinions. A growing body of work seeks to align LLMs with particular human population groups, with the goal of simulating their opinions or behaviors, by steering LLM generations towards those groups Aher et al. (2023); Sun et al. (2024); Dillion et al. (2023). One common method to do so is persona prompting, which refers to defining a persona with demographic (sociodemographic prompting) or description variables within a prompt, with the goal of modeling the behavior of that group using an LLM Hu and Collier (2024); Beck et al. (2024).
@@
Recent work in NLP has sought to evaluate sociodemographic (SD) prompting by studying its performance on large subjective surveys with thousands of human responses Santurkar et al. (2023); Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). However, evaluation of SD prompting has typically suffered from a key shortcoming: many studies assume a single majority answer to each question per sociodemographic group in modeling and/or evaluation Hwang et al. (2023); Hu and Collier (2024); Sun et al. (2025a); Mukherjee et al. (2024); Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). The fallacy in such a framing is clear: because members of a given group do not all share the same beliefs, accurate modeling and evaluation must incorporate intra-group disagreement. Prompting for a distribution, rather than a representative response, could potentially also make a model less susceptible to caricature and stereotyping Cheng et al. (2023a); Wang et al. (2025); Cheng et al. (2023b), better aligning with the goals of pluralistic alignment Sorensen et al. (2024). Prior studies that do model distributional beliefs tend to evaluate only a single method of extracting an LLM distribution Santurkar et al. (2023); Sun et al. (2024), or evaluate LLM distributions in an SD-agnostic manner Beck et al. (2024).
@@
In this work, we evaluate SD prompting in a distributional manner across three survey datasets using a variety of methods to extract distributions from LLMs. We focus on methods that broadly support use with open-source, open-model, and API-access LLMs. We apply supervised calibration (described in Section 3.6) to better align (see Section 3.4) LLM-generated distributions with human response distributions, with the intuition that LLM distributions might be directionally correct, but simply uncalibrated; e.g., LLM distributions might exaggerate differences between groups.
@@
Concretely, we explore distributional opinion alignment via four research questions: (RQ 1) Does SD prompting generate distributions that are more aligned with human opinion? (RQ 2) Can LLM-generated distributions be calibrated to better align with humans using supervised data? (RQ 3) Can LLM distributions be more easily aligned with some SD groups? (RQ 4) How many supervised training examples does calibration need?
@@
We evaluate the alignment of LLM-generated distributions to human opinion across 3 distribution elicitation methods and 15 models of varying degrees of openness (open-source, open-model, API-access only), size, pre-training data, and post-training methods. We first show that baseline SD prompting techniques do not lead to consistent improvements in alignment across settings (RQ1). In contrast, our approach to supervised calibration of LLM-generated distributions improves alignment by 16.3% on average across prompt methods, elicitation methods, datasets, and LLMs (RQ2), though it improves for some SD groups and degrades for others (RQ3). Moreover, as few as 1-10 gold examples can be used to calibrate distributions (RQ4).
Overall, we find that applying calibration to LLM distributions reduces the importance of model and elicitation method choice by reliably improving alignment. However, we emphasize caution given that alignment success varies across SD groups and evaluating with SDs risks essentializing demographic categories.
@@ 2 Related Work
As we seek to do in our work, many methods have been proposed to better align LLMs with human values and opinions Ma et al. (2024). Persona prompting—defining a persona with description variables within a prompt to simulate or model groups—is one such method Hu and Collier (2024). One form, sociodemographic (SD) prompting, defines the persona with SD variables Beck et al. (2024). Other methods include: using knowledge of a user’s previous opinions to model their future opinions Hwang et al. (2023); Gordon et al. (2022), using psychological scaffolds to improve model rationalization Joshi et al. (2025), self-alignment via the simulation of social scenes Pang et al. (2024), extracting persona vectors to mitigate certain character traits Chen et al. (2025), creating more descriptive value profiles Sorensen et al. (2025), using in-context learning Choenni and Shutova (2025), and fine-tuning LLMs followed by correction with a small amount of human data Krsteski et al. (2025). Evaluation frameworks have also been proposed for assessing alignment and social simulation Kang et al. (2025); Shi et al. (2025); Lutz et al. (2025).
@@
However, several works report inadequacies in these methods in accurately aligning models with humans, such as LLMs being unreliable as personalized annotators Dong et al. (2024), unable to use different languages to steer generation towards groups Kwok et al. (2024), biased towards some groups Sun et al. (2025b), lacking stability and steerability under prompt variations Khan et al. (2025), and lacking the reasoning depth to simulate humans in social experiments Gao et al. (2025).
Prior work evaluating SD prompting has also reported mixed results in accurately simulating social groups Beck et al. (2024); Santurkar et al. (2023); Hu and Collier (2024); Kaiser et al. (2025). While some works show that SD prompting can improve performance, results vary greatly by prompt, model, and task Beck et al. (2024); Mukherjee et al. (2024); Bisbee et al. (2024). Other work warns against risks of misportrayal, othering, and exoticization of identities Cheng et al. (2023a); Wang et al. (2025); Cheng et al. (2023b); Kitadai et al. (2024) and have shown biases in language models that may be perpetuated with SD prompting Nadeem et al. (2021); Gallegos et al. (2024); Gupta et al. (2024).
@@
Our work follows a body of work that has advocated for a shift to pluralistic alignment due to the ecological fallacy of assuming all members of a group share the same values, beliefs, or opinions Gordon et al. (2022); Sorensen et al. (2024); Feng et al. (2024); Kirk et al. (2024); Xie et al. (2025). Similar to our work, prior works have studied aligning LLM distributions with human response distributions on subjective questions. However, some of this work focuses on the systematic biases with LLM-generated distributions Dominguez-Olmedo et al. (2024) or on fine-tuning approaches Cao et al. (2025). Santurkar et al. (2023) evaluate alignment with human opinion via a large US opinion poll. We draw inspiration from this work, though we evaluate multiple methods of extracting distributions from LLMs across multiple datasets. Closest to our work, Meister et al. (2025) also benchmarks the distributional alignment of LLMs with various distribution elicitation methods and several datasets. However, we introduce supervised calibration as a method for improving distributional alignment.
@@ 3.1 Datasets
We consider three survey datasets: Welcome Global Monitor 2018 (WGM) Wellcome Trust and The Gallup Organization Ltd (2019), OpinionQA (OQA) Santurkar et al. (2023), and the World Values Survey (WVS) Haerpfer et al. (2020). OQA and WVS have been used to study alignment in prior work Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). These three datasets cover diverse subjective topics, including global perceptions of science and public health (WGM), US public opinion (OQA), and various moral opinions and values (WVS). We include all WGM ordinal questions and two questions per category from OQA and WVS, yielding 92 questions across WGM (14), OQA (38), and WVS (40) and a total of 4,500 human response distributions over all datasets, questions, and demographics.
@@ 3.2 Distribution elicitation
From ground truth human response data, we infer reference probability distributions over answer choices for each question by relative frequency, specific to each SD group. To generate distributions via LLM, we apply methods inspired from model confidence literature. We investigate techniques that work most broadly, supporting use with open-source, open-model, and API-access LLMs. That said, Appendix D.2 reports a smaller scale study with log probabilities for comparison. Appendix I shows prompts used for all methods.
Verbalized directly prompts the model to output a probability distribution Geng et al. (2024); Tian et al. (2023). For example, given the question How much do you trust vaccines?, with answer choices a lot, somewhat, not much, and not at all, we prompt the model to output a distribution over all answer choices, such as [0.7, 0.2, 0.05, 0.05]. We sample $n=3$ times and average. We re-normalize any improper distributions and discard invalid outputs.
Self-random prompts the model to output a single answer choice, samples $n=5$ times with temperature 0.7 Xiong et al. (2024), then creates a distribution over the $n$ responses. For example, given the question How much do you trust vaccines?, with answer choices a lot, somewhat, not much, and not at all, we prompt the model to output a single number that correlates with an answer choice (e.g., 4 for a lot or 1 for not at all).
Paraphrase prompts the model to output a single answer choice but uses different paraphrases of the prompt instead of sampling responses for the same prompt Xiong et al. (2024). We use $n=5$ paraphrases for each prompt.
@@ 3.3 Prompts
We use two categories of prompts, with three variations. For each, we prompt with the 3 different elicitation methods. See Appendix I for the prompts.
@@
Prompt variations. We study three prompt variations, standard prompts, few shot prompts (Section C), and chain of thought (CoT) prompts Kojima et al. (2022); Wei et al. (2022) (Section B). See Appendix Table 18 for the variations.
@@ 3.4 Metric: Opinion Alignment
where $WD$ refers to Wasserstein distance, or earth-mover’s distance, and N1N-1 is a normalization factor. Santurkar et al. (2023) note this metric accounts for the ordinal nature of the survey questions, as opposed to other distribution divergence metrics like Kullback-Liebler or Jensen-Shannon. The opinion alignment metric ranges in $[0,1]$, but we show the value as a percentage in our results.
@@ 3.5 Models
We evaluate 15 models ranging in openness (open-source, open-weight, black box), size, modality, and post-training. Model families include Claude, Llama, Mistral, OLMo-2, and Qwen. Our main results report on only the most powerful model in each family. Others are reported in Appendix D.1.
@@
Table 1: Opinion alignment before and after calibration for each dataset, LLM, and elicitation method. Each pair of columns compares the base-generated or SD-generated distributions to the calibrated distributions ($C$) for each elicitation method: paraphrase ($P$), self-random ($S$), and verbalized ($V$). Bolded values are significant between each pair (calculated via paired t-test and Bonferroni correction, see Appendix G). The best performance for each model is highlighted in yellow. The best averages across models and datasets are highlighted in blue. Results for all LLMs are shown in Appendix Table 7. Adding SD information does not consistently improve alignment, but calibration more consistently improves alignment on average and in most settings.
@@ 3.6 Calibration
We apply supervised regression to transform LLM-generated distributions to more closely align them with the ground truth human distributions. For example, for the question, How much do you trust doctors and nurses? A lot, some, not much, or not at all?, the LLM might generate a distribution such as $[0.6,0.2,0.2,0]$, while the human response distribution might be $[0.25,0.5,0.1,0.15]$. We seek to learn a transformation from the LLM-generated distributions to the human distributions that might be applied to all LLM-generated distributions. More formally, for a regression model $R$ and LLM distribution $\mathbf{D}$ parameterized by the values for each answer choice $D_{a},D_{b},D_{c},...,D_{k}]$ (where $k$ is the number of answer choices), we learn a regression such that each value is transformed using supervision from ground truth values for each answer choice $G_{a},G_{b},G_{c},...,G_{k}]$. We then re-normalize the transformed distribution.
@@
Regression model. We learn a regression model for each dataset-LLM-elicitation-prompt setting (270 regression models for 3 datasets, 15 models, 3 elicitation methods, and 2 prompt categories). We perform grid search for regressor model selection and corresponding hyperparameters (see Appendix H for details). We use the scikit-learn Pedregosa et al. (2011) implementation for each regressor.
Training. The total number of examples for each regression model is $|C|$ x $|Q|$ x $|A|$, where $C$ is the class within each sociodemographic group (e.g., female and male for sex), $Q$ is the question (e.g., how much do you trust doctors?), and $A$ is the set of answer choices per question. These values differ for each setting (see Appendix Table 15). We will refer to full examples as the full distributions not split by answer choices (total of $|C|$ x $|Q|$). We were unable to extract distributions for all questions from some of the smaller LLMs (see Appendix Table 13), affecting the number of training examples. We split the LLM distributions into train, development, and test sets by questions (80%-10%-10%). All of our reported results use the test set.
For minimal supervision (Section 6), we train on a random sample of $n$ full examples ($n$ = [1, 5, 10, 50, 100, 200, 500, 1000, 1200]) from the training set. We repeat this procedure 10 times for different random splits and average.
Evaluation. We optimize the regression models with Mean Squared Error (MSE). We re-normalize output values to sum to 1 (for a proper probability distribution) before calculating alignment with ground truth distributions. For example, if the output of $R$ for the distribution $[0.6,0.2,0.2,0]$ is $[0.3,0.7,0.1,0.1]$, we would normalize these values to $[0.25,0.59,0.08,0.08]$.